OpenStudy (anonymous):

Use substitution to determine if x = 12 is the solution to the equation 5x – 33 = 15 + x

OpenStudy (anonymous):

yes no

OpenStudy (anonymous):

5(12)-33 = 15 + 12 60-33 = 27 27 = 27 Yes.